OpenStudy (anonymous):

Factor the trinomial r^2-11r+24

OpenStudy (anonymous):

(r-3)(r-8)

OpenStudy (anonymous):

Thank you

hero (hero):

???

hero (hero):

I thought I showed you the steps

hero (hero):

You should be checking answers at this point

hero (hero):

By the way you check answers by FOILING the factored form back out again

OpenStudy (anonymous):

right I have a about 15 of these

OpenStudy (anonymous):

working on some of them now with the concept you showed me Hero it is working

OpenStudy (anonymous):

do not need help for the moment

OpenStudy (anonymous):

If you ever find yourself really stuck you can always use the following formula to factor quadratics. This is based on the quadratic formula, which, if you haven't learned/memorized already, will have to at some point anyway. :P (so, you might as well!) \[ax^2+bx+c=\displaystyle a\left(x+\frac{b+\sqrt{b^2-4ac}}{2a}\right)\left(x+\frac{b-\sqrt{b^2-4ac}}{2a}\right)\]As an exercise, you should distribute the terms on the right-hand side to check that it's really the proper factorization. If you have't seen it before, for your own edification, the Quadratic Formula is below, which gives the solutions to the equation \(ax^2+bx+c=0\). It's an equation I learned in 9th grade and it comes up a lot.\[\displaystyle x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]If you get a negative number as \(b^2-4ac\) (under the square root), that means that your equation is not factorable unless you use imaginary numbers (which, for most middle school/high school classes, means it's not factorable). I know these formulas look daunting at first, but trust me, they're very useful, and I encourage you to get some familiarity with them.

OpenStudy (saifoo.khan):

(i wil give u a medal for that) ^

OpenStudy (anonymous):

Thank you! :-)

OpenStudy (saifoo.khan):

@yakey, how do u write b^2-4ac in that same line?

OpenStudy (anonymous):

saifoo.khan - Surround with \( brackets instead of the \[ ones. The latter make "display style" equations while this other one does specifically "inline" equations.

OpenStudy (saifoo.khan):

can't do it. :/

OpenStudy (anonymous):

20b^2+11b-3 =(20b+1)(b-3)

OpenStudy (anonymous):

You need to close them too with \) just like the other ones. I couldn't put both the open and close in the same post because it thought I was doing math in between.

OpenStudy (saifoo.khan):

\[(4 b+3) (5 b-1) \]

OpenStudy (saifoo.khan):

how? answer = \( Kewl\)

OpenStudy (anonymous):

@Marv06 - That factorization isn't quite correct.

OpenStudy (anonymous):

i just want to know how u got the 4

OpenStudy (anonymous):

i'm getting dizzy lol

OpenStudy (anonymous):

4 and 5 are factors of 20 that happen to produce nice whole numbers when you factorize.

OpenStudy (anonymous):

oh ok

hero (hero):

By the way you check answers by FOILING the factored form back out again

hero (hero):

Marv, remember when I was telling you that once you put a number in front of the first term, everything changes? Well, that's what happened and you didn't give me a chance to show you

OpenStudy (anonymous):

I apologize

OpenStudy (anonymous):

I'm still working on a few now on the ones I know so far

OpenStudy (anonymous):

28s^2-37s-21

hero (hero):

Yes, I need to go over that one with you

hero (hero):

Come back to skype.

hero (hero):

And bring a notepad

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me take a smoke brake my nerves is getting bad lol

OpenStudy (anonymous):

break

hero (hero):

smoke break?

hero (hero):

really?

OpenStudy (anonymous):

ok back

hero (hero):

I'm already on

OpenStudy (anonymous):

HA!